OpenStudy (anonymous):

4/a^2+6a+9 + 9/a^2-9

OpenStudy (anonymous):

that is not correct

OpenStudy (anonymous):

yes i know a wrote 4a^2 instead of 4/a^2 - ill try again !!

OpenStudy (anonymous):

is it 4 / (a^2 + 6a + 9) + 9/(a^2 -9) ?

OpenStudy (anonymous):

u have to find the lcd

OpenStudy (anonymous):

in which case its : its 4/(x+3)(x+3) + 9 (x+3)(x-3) LCD = (x+3)^2 * (x-3)

OpenStudy (anonymous):

= 4(x-3) + 9(x+3)/ (x+3)^2 * (x-3) = 13x + 15 / (x+3)^2 * (x-3)

OpenStudy (anonymous):

i believe that right though I'm making a few silly mistakes lately